OpenStudy (anonymous):

the length of one leg of a right triangle is 9ft. The length of the hypotenuse is 3ft longer than the other leg. Find the length of the hypotenuse and the other leg.

OpenStudy (anonymous):

\[n^2 + 9^2 = (n+3)^2\]

OpenStudy (anonymous):

which gives me?

OpenStudy (anonymous):

The answer that you will have to figure out! :D

OpenStudy (anonymous):

im looking for the length of the hypotenuse and the other leg and this is my first time doing this. Can you please walk me through this once?

OpenStudy (anonymous):

Ok. Do you remember the Pythagorean theorem?

OpenStudy (anonymous):

a^2+b^2=c^2 yes

OpenStudy (anonymous):

Ok. That's the basis of what I wrote above. Since one leg was given, and the hypotenuse was expressed in terms of the other leg.

OpenStudy (anonymous):

Do you know how to expand \((n+3)^{2}\)?

OpenStudy (anonymous):

not really :/

OpenStudy (anonymous):

There is a formula but I never remember it. You can just FOIL, though.

OpenStudy (anonymous):

n^2+3^2?

OpenStudy (anonymous):

You forgot the middle terms. \((a + b)(c + d) = ab + ad + bc + bd\)

OpenStudy (anonymous):

ah ok.

OpenStudy (anonymous):

Er... ac + ad + bc + bd... sorry.

OpenStudy (anonymous):

no worries im listening

OpenStudy (anonymous):

So \((n + 3)(n + 3) = n^2 + 3n + 3n + 9\)

OpenStudy (anonymous):

So far we have \(n^2 + 81 = n^2 + 6n + 9\)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Good news, we don't have to solve a quadratic equation, because we can just subtract the \(n^2\), it should be easy to solve then.

OpenStudy (anonymous):

so the length of the hypotenuse is?

OpenStudy (anonymous):

Slow down! I'm walking you through it. :P

OpenStudy (anonymous):

alright sorry lol

OpenStudy (anonymous):

continue :)

OpenStudy (anonymous):

\[6n + 9 = 81\]\[6n = 72\]\[n = 12\]

OpenStudy (anonymous):

So now you tell me the length of the hypotenuse!

OpenStudy (anonymous):

then if the hypotenuse is 12 then the other leg is 3ft longer so its 15?

OpenStudy (anonymous):

You got that backwards but yes. The hypotenuse is the one that is 3ft longer.

OpenStudy (anonymous):

got it now i remember the hypotenuse leg was always longer than the others

OpenStudy (anonymous):

I appreciate your time and your help

OpenStudy (anonymous):

:D (Give me a medal then :P )

OpenStudy (anonymous):

I wish i could lol

OpenStudy (anonymous):

Just click the "best response" thingy next to any of my posts.

OpenStudy (anonymous):

however your actions wont go unnoticed

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

Good luck with your future mathematic endeavours

OpenStudy (anonymous):

thank you, I hate to bother you but I have one last math equation thats bothering me and ill leave you alone for the rest of tonight

OpenStudy (anonymous):

If it's a new question it's better to open a new topic. But I'll go there and help you.

OpenStudy (anonymous):

a room has a rectangular space of 230^2 for a square dining room and a 13ft wide kitchen. find both the width of the square dining room nd the length of the entire rectangular space

OpenStudy (anonymous):

I'm trying to figure out how that even makes sense.

OpenStudy (anonymous):

Ok, I think I get it.

OpenStudy (anonymous):

the length of the kitchen is 13ft and the area of the dining room is 230 square feet....its kind of hard to explain

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the dining room is 230 square feet and the width of the kitchen is 13ft

OpenStudy (anonymous):